# On Discounted Infinite-Time Mean Field Games

### v4

@@ 6 Viscosity solution to distribution-dependent elliptic PDE
Here $H(x,\mu,y)$ is defined in (5.1) and the above equations are the same as (1.5) and (1.6).
We define $\mathcal{V}(x,\mu)\triangleq Y^{x,\xi}_{0}$ and attempt to prove that it satisfy the following elliptic PDE:
@@ 6.1 Yamada-Watanabe theorem for infinite-time FBSDEs
We know that under Assumption 2.1, the above FBSDE has unique solution $(X_{t},Y_{t},Z_{t})$ in $L_{r}^{2}$ with initial $X_{0}\in\mathbb{L}^{2}({\cal F}_{0})$. We then give the definitions of strong and weak uniqueness.
@@
We say that the strong uniqueness holds for FBSDE (6.4) if on any admissible set-up $(\Omega,{\cal F},\mathbb{P},\mathbb{F})$ with input $(X_{0},B)$, for any two $\mathbb{F}$-progressively measurable $L^{2}_{r}$ three-tuples
@@
For any two set-ups $(\Omega^{1},\mathcal{F}^{1},\mathbb{P}^{1},\mathbb{F}^{1})$ and $(\Omega^{2},\mathcal{F}^{2},\mathbb{P}^{2},\mathbb{F}^{2})$ with inputs $(X_{0}^{1},B^{1})$ and $(X_{0}^{2},B^{2})$, $X_{0}^{1},X_{0}^{2}$ have the same law on $\mathbb{R}$. We say the weak uniqueness holds for FBSDE (6.4) if the $L^{2}_{r}$ solutions $(X_{t}^{1},Y_{t}^{1},\int_{0}^{t}Z^{1}_{s}\mathrm{d}s)_{t\geq 0}$ and $(X_{t}^{2},Y_{t}^{2},\int_{0}^{t}Z^{2}_{s}\mathrm{d}s)_{t\geq 0}$ on corresponding set-ups have the same distribution.
We use the same scheme as the one developed by Yamada and Watanabe to prove that path-wise uniqueness of solutions of FBSDE’s implies uniqueness in the sense of probability law.
@@
Assume that on $(\Omega,\mathcal{F},\mathbb{P},\mathbb{F})$ with input $(X_{0},B_{t})_{t\geq 0}$, the FBSDE has a unique strong solution $(X_{t},Y_{t},Z_{t})$. Then the law of $(X_{0},B_{t},X_{t},Y_{t},\int_{0}^{t}Z_{s}\mathrm{d}s)$ only depends on $\mathcal{L}(X_{0})$.
Proof.  At first we denote by $C([0,\infty);\mathbb{R})$ the space of continuous $\mathbb{R}$-valued functions on $0,\infty)$ equipped with the metric of uniform convergence on compacts:
@@
Next, we need a enlarged space $(\Omega_{\text{total}},\mathcal{G},Q)$ to support all processes. Define
@@
Let us now endow $(\Omega_{\text{total}},{\cal G},Q)$ with the filtration $\mathbb{G}$, where $\mathbb{G}=\{{\cal G}_{t}\}_{t\geq 0}$ is the complete and right-continuous augumentation under $Q$ of the canonical filtration
@@
on $\Omega_{\text{total}}$. It’s easy to see that $\xi$ is ${\cal G}_{0}$-measurable and that $(w_{s},x^{1}_{s},y^{1}_{s},z^{1}_{s},x^{2}_{s},y^{2}_{s},z^{2}_{s})_{t\geq 0}$ is $\{{\cal G}_{t}\}_{t\geq 0}$-progressively measurable. Moreover, for $i\in\{1,2\}$,
@@
Actually, we just have to prove that $(w_{t})_{t\geq 0}$ is a $\{{\cal G}_{t}\}_{t\geq 0}$ Brownian motion: we follow the proof given in ([6], Remark 1.6). Let us firstly define
@@
At last, applying the same procedure on $(z^{i}_{t})_{t\geq 0},i\in\{1,2\}$ in ([5], Volume II, Lemma 1.27), we obtain that Qa.s.Q-a.s.: for all $0\leq t\leq T,i\in\{1,2\}$,
@@
6.2 Regularity of the FBSDE solutions
@@
For FBSDEs (6.1) and (6.2), we define $\mathcal{V}(x,\mu)\triangleq Y^{x,\xi}_{0}$ for some initial state $\xi$ with distribution $\mu$. Assuming all assumptions in Theorem 4.4 are satisfied, we have, for any $x,x^{\prime}\in\mathbb{R}$ and $\mu,\mu^{\prime}\in\mathcal{P}_{2}$, there exists a constant $C>0$, such that
@@
$|\mathcal{V}(x_{1},\mu_{1})-\mathcal{V}(x_{2},\mu_{2})|^{2}\leq C_{4}\left(|x_{1}-x_{2}|^{2}+\mathcal{W}_{2}^{2}(\mu_{1},\mu_{2})\right),$ (6.35)
then we get the desired result.
6.3 Connection with distribution-dependent elliptic PDE
Now let us give the definition of a viscosity solution for PDE (6.3). We rewrite the PDE as follows:
where
$\begin{split}(\mathcal{L}\Phi)[\Psi](x,\mu)\triangleq&\partial_{y}H(x,\mu,\Psi)\cdot\partial_{x}\Phi(x,\mu)+\frac{1}{2}\partial_{xx}\Phi(x,\mu)\\ &+\tilde{\mathbb{E}}\left[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}\Phi(x,\mu,\tilde{\xi})+\partial_{\mu}\Phi(x,\mu,\tilde{\xi})\partial_{y}H(\tilde{\xi},\mu,\Psi(\tilde{\xi},\mu))\right],\end{split}$ (6.37)
and
Definition 6.6
[…]
and
[…]
Definition 6.7
(respectively
We now assert that
@@ 6.3 Connection with distribution-dependent elliptic PDE
Assume that both FBSDEs (6.1) and (6.2) admit unique solutions in $L^{2}_{r}$, $Y_{0}^{x,\xi}$ depends solely on the distribution of $\xi$. Then we define the function $\mathcal{V}(x,\mu)\triangleq Y_{0}^{x,\xi}$ with $\mathcal{L}_{\xi}=\mu$, and assume $\mathcal{V}$ is jointly continuous on $\mathbb{R}\times\mathcal{P}_{2}$. The function $\mathcal{V}(x,\mu)$ is a viscosity solution of PDE (6.36).
Proof.  We only show that $\mathcal{V}$ is a viscosity subsolution of PDE (6.36). A similar argument will show that it is also a viscosity supersolution of (6.36).
Due to the uniqueness for solutions of FBSDE, it is not hard to see that for any $t\geq 0$, $\mathcal{V}(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}})=Y_{t}^{x,\xi}$. Let $\Psi$ be a test function, and $(x^{0},\mu^{0})\in\mathbb{R}\times\mathcal{P}_{2}$ be a local maximum of $\mathcal{U}-\Psi$. We assume without loss of generality that $\mathcal{V}(x^{0},\mu^{0})=\Psi(x^{0},\mu^{0})$. We suppose that
@@
It follows from the above that there exists an open subset $O\subset\mathbb{R}\times\mathcal{P}_{2}$ that contains $(x^{0},\mu^{0})$, such that for all $(x,\mu)\in O$,
@@
Taking a initial state $\xi^{0}\in\mathbb{L}^{2}(\mathcal{F}_{0};\mu^{0})$, we consider the processes $(X_{t}^{\xi^{0}},Y_{t}^{\xi^{0}},Z_{t}^{\xi^{0}})$ and $(X_{t}^{x^{0},\xi^{0}},Y_{t}^{x^{0},\xi^{0}},Z_{t}^{x^{0},\xi^{0}})$ which are solutions to FBSDEs (6.1) and (6.2). For some $T>0$, let $\tau$ denote the stopping time

### v5

@@ 6 Viscosity solution to distribution-dependent elliptic PDE
Here, $H(x,\mu,y)$ is defined in (5.1), and the above equations are the same as (1.6) and (1.7).
We define $\mathcal{V}(x,\mu)\triangleq Y^{x,\xi}_{0}$ and attempt to prove that it satisfies the following elliptic PDE in the viscosity sense:
@@
Now, let us give the definition of a viscosity solution for PDE (6.3). We rewrite the PDE as follows:
where
$\begin{split}(\mathcal{L}\Phi)[\Psi](x,\mu)\triangleq&\partial_{y}H(x,\mu,\Psi(x,\mu))\cdot\partial_{x}\Phi(x,\mu)+\frac{1}{2}\partial_{xx}\Phi(x,\mu)\\ &+\tilde{\mathbb{E}}\left[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}\Phi(x,\mu,\tilde{\xi})+\partial_{\mu}\Phi(x,\mu,\tilde{\xi})\partial_{y}H(\tilde{\xi},\mu,\Psi(\tilde{\xi},\mu))\right],\end{split}$ (6.5)
and
Definition 6.1
[…]
and
[…]
Definition 6.2
(respectively,
@@
We will next establish them sequentially in this section and ultimately prove that $\mathcal{V}$ is a viscosity solution to equation (6.3)
@@ 6.1 Yamada-Watanabe theorem for infinite-time FBSDEs
We know that under Assumption 2.1, the above FBSDE has a unique solution $(X_{t},Y_{t},Z_{t})$ in $L_{r}^{2}$ with initial $X_{0}\in\mathbb{L}^{2}({\cal F}_{0})$. We then give the definitions of strong and weak uniqueness.
@@
We say that the strong uniqueness holds for FBSDE (6.11) if on any admissible setup $(\Omega,{\cal F},\mathbb{P},\mathbb{F})$ with input $(X_{0},B)$, for any two $\mathbb{F}$-progressively measurable $L^{2}_{r}$ three-tuples
@@
For any two set-ups $(\Omega^{1},\mathcal{F}^{1},\mathbb{P}^{1},\mathbb{F}^{1})$ and $(\Omega^{2},\mathcal{F}^{2},\mathbb{P}^{2},\mathbb{F}^{2})$ with inputs $(X_{0}^{1},B^{1})$ and $(X_{0}^{2},B^{2})$, $X_{0}^{1},X_{0}^{2}$ have the same law on $\mathbb{R}$. We say the weak uniqueness holds for FBSDE (6.11) if for the $L^{2}_{r}$ solutions $(X_{t}^{1},Y_{t}^{1},Z^{1}_{t})_{t\geq 0}$ and $(X_{t}^{2},Y_{t}^{2},Z^{2}_{t})_{t\geq 0}$ on corresponding set-ups, the processes $(X_{t}^{1},Y_{t}^{1},\int_{0}^{t}Z^{1}_{s}\mathrm{d}s)_{t\geq 0}$ and $(X_{t}^{2},Y_{t}^{2},\int_{0}^{t}Z^{2}_{s}\mathrm{d}s)_{t\geq 0}$ have the same distribution.
We use the same scheme as the one developed by Yamada and Watanabe to prove that pathwise uniqueness of solutions of FBSDE’s implies uniqueness in the sense of probability law.
@@
Assume that on $(\Omega,\mathcal{F},\mathbb{P},\mathbb{F})$ with input $(X_{0},B_{t})_{t\geq 0}$, the FBSDE has a unique strong solution $(X_{t},Y_{t},Z_{t})$. Then, the law of $(X_{0},B_{t},X_{t},Y_{t},\int_{0}^{t}Z_{s}\mathrm{d}s)$ only depends on $\mathcal{L}(X_{0})$.
Proof.  First, we denote by $C([0,\infty);\mathbb{R})$ the space of continuous $\mathbb{R}$-valued functions on $0,\infty)$ equipped with the metric of uniform convergence on compacts:
@@
Next, we need an enlarged space $(\Omega_{\text{total}},\mathcal{G},Q)$ to support all processes. Define
@@
Let us now endow $(\Omega_{\text{total}},{\cal G},Q)$ with the filtration $\mathbb{G}$, where $\mathbb{G}=\{{\cal G}_{t}\}_{t\geq 0}$ is the complete and right-continuous augmentation under $Q$ of the canonical filtration
@@
on $\Omega_{\text{total}}$. It is easy to see that $\xi$ is ${\cal G}_{0}$-measurable and that $(w_{s},x^{1}_{s},y^{1}_{s},z^{1}_{s},x^{2}_{s},y^{2}_{s},z^{2}_{s})_{t\geq 0}$ is $\{{\cal G}_{t}\}_{t\geq 0}$-progressively measurable. Moreover, for $i\in\{1,2\}$,
@@
Actually, we just have to prove that $(w_{t})_{t\geq 0}$ is a $\{{\cal G}_{t}\}_{t\geq 0}$ Brownian motion: we follow the proof given in [6] (Remark 1.6). Let us firstly define
@@
At last, applying the same procedure to $(z^{i}_{t})_{t\geq 0},i\in\{1,2\}$ in [5] (Volume II, Lemma 1.27), we obtain that Qa.s.Q-a.s.: for all $0\leq t\leq T,i\in\{1,2\}$,
@@
6.2 Connection with distribution-dependent elliptic PDE
For FBSDEs (6.1) and (6.2), we define $\mathcal{V}(x,\mu)\triangleq Y^{x,\xi}_{0}$ for some initial state $\xi$ with distribution $\mu$. We first present the continuity result for $\mathcal{V}$.
@@ 6.2 Connection with distribution-dependent elliptic PDE
Assuming all assumptions in Theorem 4.4 are satisfied, we have, for any $x,x^{\prime}\in\mathbb{R}$ and $\mu,\mu^{\prime}\in\mathcal{P}_{2}$, there exists a constant $C>0$ such that
@@
$|\mathcal{V}(x_{1},\mu_{1})-\mathcal{V}(x_{2},\mu_{2})|^{2}\leq C_{4}\left(|x_{1}-x_{2}|^{2}+\mathcal{W}_{2}^{2}(\mu_{1},\mu_{2})\right).$ (6.42)
Then, we get the desired result.
We now assert the following theorem.
@@
Considering all assumptions in Theorem 4.4 are satisfied, the function $\mathcal{V}(x,\mu)$ is a viscosity solution of PDE (6.4).
Proof.  We only show that $\mathcal{V}$ is a viscosity subsolution of PDE (6.4). A similar argument will show that it is also a viscosity supersolution of PDE (6.4).
Due to the uniqueness of solutions of FBSDEs, it is not hard to see that for any $t\geq 0$, $\mathcal{V}(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}})=Y_{t}^{x,\xi}$. Let $\Psi$ be a test function and $(x^{0},\mu^{0})\in\mathbb{R}\times\mathcal{P}_{2}$ be a local maximum of $\mathcal{U}-\Psi$. We assume without loss of generality that $\mathcal{V}(x^{0},\mu^{0})=\Psi(x^{0},\mu^{0})$. We suppose that
@@
It follows from the above that there exists an open subset $O\subset\mathbb{R}\times\mathcal{P}_{2}$ that contains $(x^{0},\mu^{0})$ such that for all $(x,\mu)\in O$,
@@
Taking an initial state $\xi^{0}\in\mathbb{L}^{2}(\mathcal{F}_{0};\mu^{0})$, we consider the processes $(X_{t}^{\xi^{0}},Y_{t}^{\xi^{0}},Z_{t}^{\xi^{0}})$ and $(X_{t}^{x^{0},\xi^{0}},Y_{t}^{x^{0},\xi^{0}},Z_{t}^{x^{0},\xi^{0}})$, which are solutions to FBSDEs (6.1) and (6.2), respectively. For some $T>0$, let $\tau$ denote the stopping time
@@
Acknowledgements
Song Y. is financially supported by National Key R&D Program of China (No. 2024YFA1013503 & No. 2020YFA0712700) and the National Natural Science Foundation of China (No. 12431017).